OpenStudy (anonymous):

Calculate the mass of 162 L of Chlorine gas, measured at STP.

OpenStudy (abb0t):

@ STP, use 22.7 mol/L

OpenStudy (abb0t):

once you get mol, convert to mass, using molar mass (look at your periodic table) 162 L of Cl\(\sf _{(g)}\) \(\times\) \(\sf \frac{22.7~mol}{L}\) \(\times\) molar mass of Chlorinee